# Plugging Schema Graph into Multi-Table QA: A Human-Guided Framework for Reducing LLM Reliance

### v1

@@ Abstract
Large language models (LLMs) have shown promise in table Question Answering (Table QA). However, extending these capabilities to multi-table QA remains challenging due to unreliable schema linking across complex tables. Existing methods based on semantic similarity work well only on simplified hand-crafted datasets and struggle to handle complex, real-world scenarios with numerous and diverse columns. To address this, we propose a graph-based framework that leverages human-curated relational knowledge to explicitly encode schema links and join paths. Given a natural language query, our method searches this graph to construct interpretable reasoning chains, aided by pruning and sub-path merging strategies to enhance efficiency and coherence. Experiments on both standard benchmarks and a realistic, large-scale dataset demonstrate the effectiveness of our approach. To our knowledge, this is the first multi-table QA system applied to truly complex industrial tabular data.
@@ 1 Introduction
Table Question Answering (Table QA) involves answering natural language questions using structured tabular data. This capability is increasingly critical in numerous practical domains, including healthcare Singhal et al. (2025), finance Srivastava et al. (2024), and transportation Zhang et al. (2024c), where data is commonly stored in relational databases or spreadsheet-like tables. Effective Table QA systems allow non-expert users to query structured data, thus enhancing accessibility, facilitating data-driven decision-making, and reducing the reliance on manual data manipulation or complex query languages such as SQL.
@@
Multi-table QA introduces greater complexity, as answering a question often requires reasoning over multiple related tables (middle of Figure 1). This setting poses unique challenges such as identifying relevant tables, aligning schemas, and inferring join paths—difficulties not present in single-table QA. Existing methods attempt to address these through strategies like flattening and concatenating multiple tables Pal et al. (2023); Wang et al. (2024), leveraging key constraints for semantic alignment Chen et al. (2024), or adopting multi-step retrieval pipelines Wu et al. (2025a).
However, existing multi-table QA methods have been evaluated only on simplified benchmarks. For instance, the Multi-table and Multi-hop Question Answering (MMQA) benchmark Wu et al. (2025a) contains tables with an average of just five columns—far less complex than real-world datasets like Crash Investigation Sampling System (CISS) National Highway Traffic Safety Administration (2024), where tables can have up to 207 columns. These methods typically rely on heavy model training to learn schema linking or semantic alignment, which may suffice for small benchmarks but becomes highly impractical for large-scale tables. Even state-of-the-art LLMs struggle with accurate column-level matching at this scale Zhang et al. (2024a, b). As schema linking across complex tables remains a major bottleneck, current approaches fail to generalize to realistic industrial scenarios.
@@
In this paper, we introduce SGAM (Schema Graph-Assisted Multi-table QA), a graph-based framework that integrates human-curated schema knowledge to support multi-table QA, as shown in Figure 1 (bottom). Specifically, we construct a relational graph capturing the semantic and logical dependencies among table columns. Each graph node encodes essential schema details—such as column name, associated table, and primary key status—while edges represent both intra-table semantic relationships and inter-table joinable or derivation-based links. Given a natural language query, our method searches directed paths within this graph to derive clear and interpretable reasoning chains, significantly simplifying multi-table reasoning tasks. Furthermore, we introduce efficient path pruning and sub-path merging strategies to minimize redundant joins and enhance inference coherence.
@@ 3 Method
Figure 2: Illustration of our Schema Graph Assist Multi-table QA framework. Given a complex user query and multiple tables, Step 1 retrieves relevant attributes from the tables using embedding-based semantic similarity. Step 2 uses a schema graph to automatically identify and traverse the minimal set of necessary join paths, such as GV.CASEID = AVOID.CASEID and GV.VEHNO = AVOID.VEHNO, to connect related information across tables. This graph-guided reasoning enables the system to accurately extract the target attribute (OCC.AGE) by aligning semantically related fields and resolving multi-hop dependencies. In this case, the system correctly infers the driver’s age as 88.
@@ 3.2 Graph based information extraction
To address this, we apply the following pruning procedure to each path in $\mathbb{P}$. For a path $P$, let $e^{*}$ denote the last inter-table edge from the attribute node. We truncate the path to end at $e^{*}$, retaining only the suffix sub-path $P^{\prime}\subseteq P$ that lies entirely within a single table:
@@
where $e^{*}.dst$ denotes the target node of the last inter-table edge. This ensures each retained sub-path $P^{\prime}$ stays within a single table and links a relevant attribute to its terminal key. Merging these sub-paths yields a minimal, non-redundant set of attribute-to-key chains, avoiding duplicate extraction, maintaining contextual coherence, and reducing computational cost
@@ 4 Experiments
We evaluate our graph-based method through two experiments: SQL generation on the BIRD benchmark and end-to-end QA on the real-world CISS dataset. We now detail the setup and results.
@@ 4.1 Datasets
CISS: To showcase the real-world applicability of our approach, we construct a new multi-table question answering dataset based on the Crash Investigation Sampling System (CISS) National Highway Traffic Safety Administration (2024). All tables in this dataset are domain-specific and task-relevant, with well-defined key relationships. CISS provides in-depth crash data from the United States, including detailed information on crash types, accidents’ causes, injury severity, vehicle types, and more. The dataset covers approximately 3000 to 4500 crashes annually. The recorded crash cases span from accident year 2017 to 2023. We began by converting structured tabular data into textual summaries and use GPT-4o to automatically generate question-answer pairs. Each question is manually verified for clarity, relevance, and correctness, resulting in a verified accuracy of 98%. The final dataset consists of more than 620 high-quality QA pairs covering diverse aspects of traffic accidents, including crash types, vehicle conditions, driver and pedestrian demographics, and injury outcomes. To encourage varied reasoning complexity, the dataset includes 220 one-hop, 300 two-hop, and 100 multi-hop questions, where a hop refers to a distinct reasoning step required to connect different pieces of information across attributes or tables. This CISS-based QA dataset also serves as a case study to evaluate the effectiveness of our graph-enhanced approach in complex, domain-driven applications.
@@
Numerical 78 46 - -
List 71 39 - -
Select 11 12 - -
Count 9 9 - -
@@ 4.4 Tasks and Evaluation Metrics
End-to-end evaluation (using CISS): Unlike traditional Text2SQL tasks that assume clean, well-structured data, the CISS dataset presents real-world challenges such as mixed structured and textual fields, inconsistent values, and loosely connected tables. Many attributes can only be used when combined with other attributes, and reasoning often goes beyond simple table joins. To demonstrate that a schema graph can not only guide SQL generation but also support tasks beyond SQL’s capability—such as resolving inconsistencies and modeling cross-table semantics—we conduct experiments on a real-world QA task over CISS.
@@ 4.5 Results: Graph-Aware Evaluation
Model Oracle Knowledge Execution on
@@
Distillery + GPT-4o Maamari et al. (2024) 67.21 71.83
RSL-SQL + GPT-4o Cao et al. (2024) 68.12 70.21
OpenSearch-SQL Xie et al. (2025) 69.30 72.28
AskData + GPT-4o (AT&T CDO-DSAIR) 11footnotetext: This data is collected from: https://bird-bench.github.io/ 75.36 77.14
AskData + GPT-4o (AT&T CDO-DSAIR) 11footnotetext: This data is collected from: https://bird-bench.github.io/ 65.91 (-9.45) 67.41 (-9.73)
@@ 4.6 Results: End-to-end Evaluation
To validate the applicability of our method in real-world, noisy, and semantically entangled data settings, we further evaluate it on the CISS dataset, which poses challenges beyond what SQL alone can handle.
@@ A.3 Attributes select for SQL
1. A database schema table with sheet__\__name (table), column__\__name (field), and key__\__words (keywords associated with the field).
@@
Your task is: Split the question into key semantic fragments (e.g., person name, gender, event, year, etc.). For each fragment, return the most relevant column__\__name and its corresponding sheet__\__name based on the key__\__words. Ensure that each value in the output matches the database **exactly**, including full names, suffixes (e.g., "II", "Jr."), and punctuation (e.g., commas).
@@ A.4 SQL generation
Only join tables that contribute directly to the SELECT, WHERE, or GROUP BY clauses. Avoid unnecessary joins (e.g., ‘games__\__competitor‘) if the question does not involve participation, age, or game-specific attributes. Unused joins may introduce incorrect row duplication.
@@
Path 1: full__\__name[person], id[person], person__\__id[games__\__competitor], id[games__\__competitor]
Path 2: games__\__name[games], id[games], games__\__id[games__\__competitor], id[games__\__competitor]
Path 3: age[games__\__competitor], id[games__\__competitor]
@@
‘‘‘sql SELECT games.games__\__name FROM person JOIN games__\__competitor ON person.id = games__\__competitor.person__\__id JOIN games ON games.id = games__\__competitor.games__\__id JOIN competitor__\__event ON games__\__competitor.id = competitor__\__event.competitor__\__id WHERE person.full__\__name = ’John Aalberg’ AND games__\__competitor.age = 31; ‘‘‘
@@ Appendix B Case Study
Information extraction result: In case 27187, for vehicle NO.2, the total speed is: Actual value 38, the longitudinal speed is: Actual value -37 In case 27187, for vehicle NO.1, the model is: Grand Prix

### v3

@@ Abstract
Large language models (LLMs) have shown promise in table Question Answering (Table QA). However, extending these capabilities to multi-table QA remains challenging due to unreliable schema linking across complex tables. Existing methods based on semantic similarity work well only on simplified hand-crafted datasets and struggle to handle complex, real-world scenarios with numerous and diverse columns. To address this, we propose a graph-based framework that leverages human-curated relational knowledge to explicitly encode schema links and join paths. Given a natural language query, our method searches on graph to construct interpretable reasoning chains, aided by pruning and sub-path merging strategies to enhance efficiency and coherence. Experiments on both standard benchmarks and a realistic, large-scale dataset demonstrate the effectiveness of our approach. To our knowledge, this is the first multi-table QA system applied to truly complex industrial tabular data.
@@ 1 Introduction
Table Question Answering (Table QA) involves answering natural language questions using structured tabular data. This capability is increasingly critical in numerous practical domains, including healthcare Singhal et al. (2025), finance Srivastava et al. (2024), and transportation Zhang et al. (2024c), where data is commonly stored in relational databases Wang and Yu (2025) or spreadsheet-like tables. Effective Table QA systems allow non-expert users to query structured data, thus enhancing accessibility, facilitating data-driven decision-making, and reducing the reliance on manual data manipulation or complex query languages such as SQL.
@@
Multi-table QA introduces greater complexity, as answering a question often requires reasoning over multiple related tables (middle of Figure 1). This setting poses unique challenges such as identifying relevant tables, aligning schemas, and inferring join paths, which are not present in single-table QA. Existing methods attempt to address these through flattening and concatenating multiple tables Pal et al. (2023); Wang et al. (2024), leveraging key constraints for semantic alignment Chen et al. (2024), or adopting multi-step retrieval pipelines Wu et al. (2025a).
However, existing multi-table QA methods have been evaluated only on simplified benchmarks. For instance, the Multi-table and Multi-hop Question Answering (MMQA) benchmark Wu et al. (2025a) contains tables with an average of just five columns. These are far less complex than real-world datasets like Crash Investigation Sampling System (CISS) National Highway Traffic Safety Administration (2024), where tables can have up to 207 columns. These methods typically rely on heavy model training to learn schema linking or semantic alignment, which may suffice for small benchmarks but becomes highly impractical for large-scale tables. Even state-of-the-art LLMs struggle with accurate column-level matching at this scale Zhang et al. (2024a, b). As schema linking across complex tables remains a major bottleneck, current approaches fail to generalize to realistic industrial scenarios.
@@
In this paper, we introduce SGAM (Schema Graph-Assisted Multi-table QA), a graph-based framework that integrates human-curated schema knowledge to support multi-table QA, as shown in Figure 1 (bottom). Specifically, we construct a relational graph capturing the semantic and logical dependencies among table columns. Each graph node encodes essential schema details, such as column name, associated table, and primary key status, while edges represent both intra-table semantic relationships and inter-table joinable or derivation-based links. Given a natural language query, our method searches directed paths within this graph to derive clear and interpretable reasoning chains, significantly simplifying multi-table reasoning tasks. Furthermore, we introduce efficient path pruning and sub-path merging strategies to minimize redundant joins and enhance inference coherence.
@@ 3 Method
Figure 2: Illustration of our Schema Graph Assist Multi-table QA framework. Given a complex user query and multiple tables, Step 1 retrieves relevant attributes from the tables using embedding-based semantic similarity. Step 2 uses a schema graph to automatically identify and traverse the minimal set of necessary join paths, such as GV.CASEID = AVOID.CASEID and GV.VEHNO = AVOID.VEHNO, to connect related information across tables. Step 3 uses LLM to generate a final answer based on extracted information. This graph-guided reasoning enables the system to accurately extract the target attribute (OCC.AGE) by aligning semantically related fields and resolving multi-hop dependencies. In this case, the system correctly infers the driver’s age as $8$.
@@ 3.2 Graph based information extraction
To address this, we apply the following pruning procedure to each path in $\mathbb{P}$. For a path $P$, let $e^{*}$ denote the last inter-table edge from the attribute node. We truncate the path to end at $e^{*}$, retaining only the sub-path $P^{\prime}\subseteq P$ that lies entirely within a single table:
@@
where $e^{*}.dst$ denotes the target node of the last inter-table edge. This ensures each retained sub-path $P^{\prime}$ stays within a single table and links a relevant attribute to its terminal key. Merging these $P^{\prime}$ yields a minimal, non-redundant set of attribute-to-key chains, avoiding duplicate extraction, maintaining contextual coherence, and reducing computational cost
@@ 4 Experiments
We evaluate our graph-based method through two experiments: SQL generation on the BIRD benchmark and end-to-end QA on the real-world CISS dataset. We now detail the setup and results  [^1] https://github.com/hiXixi66/SGAM-Multi-table-QA.
@@ 4.1 Datasets
CISS: To showcase the real-world applicability of our approach, we construct a new multi-table question answering dataset based on the CISS. All tables in this dataset are domain-specific and task-relevant, with well-defined key relationships. CISS provides in-depth crash data from the United States, including detailed information on crash types, accidents’ causes, injury severity, vehicle types, and more. The dataset covers approximately 3000 to 4500 crashes annually. The recorded crash cases span from accident year 2017 to 2023. We began by converting structured tabular data into textual summaries and use GPT-4o to automatically generate question-answer pairs. Each question is manually verified for clarity, relevance, and correctness, resulting in a verified accuracy of 98%. The final dataset consists of more than 620 high-quality QA pairs covering diverse aspects of traffic accidents, including crash types, vehicle conditions, driver and pedestrian demographics, and injury outcomes. To encourage varied reasoning complexity, the dataset includes 220 one-hop, 300 two-hop, and 100 multi-hop questions, where a hop refers to a distinct reasoning step required to connect different pieces of information across attributes or tables. This CISS-based QA dataset also serves as a case study to evaluate the effectiveness of our graph-enhanced approach in complex, domain-driven applications.
@@
Numerical 78 46 - 5
List 71 39 - 75
Select 11 12 - 548
Count 9 9 - 0
@@ 4.4 Tasks and Evaluation Metrics
End-to-end evaluation (using CISS): Unlike traditional Text2SQL tasks that assume clean, well-structured data, the CISS dataset presents real-world challenges such as mixed structured and textual fields, inconsistent values, and complicated connected tables. Many attributes can only be used when combined with other attributes, and reasoning often goes beyond simple table joins. To demonstrate that a schema graph can be used beyond guide SQL generation tasks, such as resolving inconsistencies and modeling cross-table semantics, we conduct experiments on a real-world QA task over CISS.
@@ 4.5 Results: Graph-Aware Evaluation
Model
Oracle
Knowledge
Execution on
@@
Distillery + GPT-4o
Maamari et al. (2024)
67.21 71.83
RSL-SQL + GPT-4o
Cao et al. (2024)
68.12 70.21
OpenSearch-SQL
Xie et al. (2025)
69.30 72.28
AskData + GPT-4o
(AT&T CDO-DSAIR)
22footnotetext: This data is collected from: https://bird-bench.github.io/
75.36 77.14
AskData + GPT-4o
(AT&T CDO-DSAIR)
22footnotetext: This data is collected from: https://bird-bench.github.io/
65.91 (-9.45) 67.41 (-9.73)
@@ 4.6 Results: End-to-end Evaluation
To validate the applicability of our method in real-world, noisy, and semantically entangled data settings, we further evaluate it on the CISS dataset.
@@ 7 Limitations
Funding sources
Part of this research was supported by the Chalmers Transport Area of Advance project TREND.
Acknowledgements
The authors would like to thank e-Commons (research infrastructure at Chalmers University of Technology) for their consultations, and extend special thanks to Nora Speicher for valuable discussions related to running the experiments. We acknowledge the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725, for awarding this project access to the Alvis cluster for computations and data handling.
@@ A.3 Attributes select for SQL
1. A database schema table with sheet_\_name (table), column_\_name (field), and key_\_words (keywords associated with the field).
@@
Your task is: Split the question into key semantic fragments (e.g., person name, gender, event, year, etc.). For each fragment, return the most relevant column_\_name and its corresponding sheet_\_name based on the key_\_words. Ensure that each value in the output matches the database **exactly**, including full names, suffixes (e.g., "II", "Jr."), and punctuation (e.g., commas).
@@ A.4 SQL generation
Only join tables that contribute directly to the SELECT, WHERE, or GROUP BY clauses. Avoid unnecessary joins (e.g., ‘games_\_competitor‘) if the question does not involve participation, age, or game-specific attributes. Unused joins may introduce incorrect row duplication.
@@
Path 1: full_\_name[person], id[person], person_\_id[games_\_competitor], id[games_\_competitor]
Path 2: games_\_name[games], id[games], games_\_id[games_\_competitor], id[games_\_competitor]
Path 3: age[games_\_competitor], id[games_\_competitor]
@@
‘‘‘sql SELECT games.games_\_name FROM person JOIN games_\_competitor ON person.id = games_\_competitor.person_\_id JOIN games ON games.id = games_\_competitor.games_\_id JOIN competitor_\_event ON games_\_competitor.id = competitor_\_event.competitor_\_id WHERE person.full_\_name = ’John Aalberg’ AND games_\_competitor.age = 31; ‘‘‘
@@ Appendix B Case Study
Information extraction result: In case 27187, for vehicle NO.2, the total speed is: Actual value 38, the longitudinal speed is: Actual value -37 In case 27187, for vehicle NO.2, the model is: Grand Prix